# MultiAiTutor: Child-Friendly Educational Multilingual Speech Generation Tutor with LLMs

### v1

@@ 4.3. Multilingual Speech Intelligence
To comprehensively assess the pronunciation accuracy of the proposed approach, we perform both objective and subjective evaluations across the three low-resource languages, as presented in Table 1. For objective evaluations, we utilize automatic speech recognition (ASR) for each language: the Tamil and Malay ASR models from (Liu et al., 2025), and the Whisper model [^1] https://huggingface.co/openai/whisper-large-v3 for Mandarin. The synthesized speech samples are transcribed using these models, and CER is calculated by comparing the transcriptions to the ground-truth text. A lower CER reflects higher pronunciation accuracy.
@@
Table 1 shows that the proposed MultiAiTutor achieves consistently better performance than the baselines across all three languages, as measured by both objective evaluations using CER and subjective evaluations by human listeners. In terms of CER, MultiAiTutor substantially reduces recognition errors compared to the baselines across three languages. These improvements reflect a higher degree of pronunciation clarity in the generated speech, making it more intelligible to ASR systems. Notably, the improvements for Tamil (from 55.08 to 83.29) and Malay (from 61.25 to 76.40) in human-rated pronunciation accuracy are substantial, highlighting the effectiveness of the proposed approach in low-resource language settings. These results confirm the capability of MultiAiTutor to advance multilingual speech generation through the integration of LLM-based modeling.
Figure 5. Comparison of Speech Quality and Naturalness Results Based on MOS Scores with 95% Confidence Intervals between SingaKids and the proposed MultiAiTutor for Malay.
@@
Speech Intelligence CosyBase MultiAiTutor
@@
Speech Intelligence SingaKids MultiAiTutor
@@
Speech Intelligence SingaKids MultiAiTutor
@@ 5. Conclusion
We propose a multilingual generative AI tutor, MultiAiTutor, designed specifically for child-friendly speech generation and educational purposes. Our approach leverages advanced LLM architectures and culturally relevant, age-appropriate training strategies to effectively support language learning in low-resource languages such as Malay, Tamil, and Singaporean-accented Mandarin. The proposed system fosters a supportive and engaging learning environment, reducing cognitive load and enhancing learner participation. Synthesized multilingual speech samples from different models are available for verification at the link  [^2] https://xiaoxue1117.github.io/icmi2025demo/.

### v3

@@ 4.3. Multilingual Speech Intelligence
To comprehensively assess the pronunciation accuracy of MultiGen, we perform both objective and subjective evaluations across the three low-resource languages, as presented in Table 1. For objective evaluations, we utilize automatic speech recognition (ASR) for each language: the Tamil and Malay ASR models we developed in SingaKids (Liu et al., 2025), and the Whisper model [^1] https://huggingface.co/openai/whisper-large-v3 for Mandarin. The synthesized speech samples are transcribed using these models, and CER is calculated by comparing the transcriptions to the ground-truth text. A lower CER reflects higher pronunciation accuracy.
@@
Table 1 shows that the proposed MultiGen achieves consistently better performance than the baselines across all three languages, as measured by both objective evaluations using CER and subjective evaluations by human listeners. In terms of CER, MultiGen substantially reduces recognition errors compared to the baselines across three languages. These improvements reflect a higher degree of pronunciation clarity in the generated speech, making it more intelligible to ASR systems. Notably, the improvements for Tamil (from 55.08 to 83.29) and Malay (from 61.25 to 76.40) in human-rated pronunciation accuracy are substantial, highlighting the effectiveness of the proposed approach in low-resource language settings. These results confirm the capability of MultiGen to advance multilingual speech generation.
Figure 5. Comparison of Speech Quality and Naturalness Results Based on MOS Scores with 95% Confidence Intervals between VITS and the proposed MultiGen for Malay.
@@
Speech Intelligence CosyBase MultiGen
@@
Speech Intelligence VITS MultiGen
@@
Speech Intelligence VITS MultiGen
@@ 5. Conclusion
We propose a multilingual generator, MultiGen, designed specifically for child-friendly speech generation. Our approach leverages advanced LLM architectures and culturally relevant, age-appropriate training strategies, and improve the text-to-speech performance in low-resource languages including Malay, Tamil, and Singaporean-accented Mandarin. The proposed approach can support a supportive and engaging communication environment with child-friendly voice, enhancing multilingual listening and speaking participation. Synthesized multilingual speech samples from different models are available for verification at the link  [^2] https://xiaoxue1117.github.io/icmi2025demo/.
6. Acknowledgment
This research is supported by A*STAR under its Japan-Singapore Joint Call: Japan Science and Technology Agency (JST) and A*STAR 2024 (R24I6IR136), and by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme (DesCartes). The educational use case is built on SingaKids (Liu et al., 2025) from A*STAR.